# DESY/PUBDB-2024-01826

### v1

@@ 5.4 Importance of Variable Classes and Individual Variables
Aiming to estimate the relative importance of a given variable class (e.g. SV variables), the classifier performance was evaluated where the given variable class was shuffled amongst all other jets, keeping the rest of the variables unchanged. The performance change concerning $b$ vs $c$, $c$ vs $s$, and $s$ vs $ud$ jet discrimination was considered with respect to the baseline, where no variable classes were shuffled. Charged jet constituent variables were found to be the most impactful variable class for all types of discrimination at a background efficiency of $\epsilon_{bkg}=10\%$. This is presumably due to charged particles being the majority of the reconstructed particles in the jets. SV variables primarily benefited $c$ vs. $s$ discrimination, with $s$ vs. $ud$ tagging particularly insensitive. Of the remaining three variable classes, V0 variables and neutral jet constituent variables were found to almost exclusively impact the performance of $s$ vs $ud$ discrimination, with little impact on both $b$ vs $c$ and $c$ vs $s$ discrimination, justifying the inclusion of V0s for identifying $s$-jets through conservation of strangeness. Moving to the high purity regime at a background efficiency of $\epsilon_{bkg}=0.1\%$, primarily the same trends were observed, with the impact of any variable type being amplified. SV variables, in particular, became hugely important to heavy flavour tagging, reaching almost equal in impact to the charged jet constituent variables, proving that the presence and properties of SVs are definitive indicators for identifying heavy flavour jets.
@@
  • Kinematic ($|p|$, $E$, $p/p_{jet.}$, $\theta$, $\Delta\theta$, …)
@@
It was found that, at a background efficiency of $10\%$, kinematic variables of charged particle constituents, including $\dfrac{E_{\text{ch.}}}{E_{\text{jet}}}$ and $\dfrac{|p_{\text{ch.}}|}{|p_{\text{jet}}|}$, were generally impactful, particularly for $c$ vs $s$ discrimination. Track variables, such as $D_{0}/\sigma_{D_{0}}$ and $Z_{0}$, were likewise impactful, though less for $b$ vs $c$ discrimination, possibly due to their redundant information after the inclusion of SVs. PID variables had little impact on $b$ vs $c$ and $c$ vs $s$ discrimination, but $K^{\pm}\text{ID}$ and photon ID were the most important for $s$ vs $ud$ discrimination, as was observed earlier. The high purity regime at a background efficiency of $0.1\%$ resulted in similar trends, though with PID variables, including $K^{\pm}\text{ID}$ and photon ID, decreasing in importance and being somewhat replaced by kinematic ones. It should be stated that the baseline $K^{\pm}\text{ID}$ scenario, as mentioned in Section 5, is deliberately pessimistic, which could account for its decrease in importance. Track variables remained impactful. The secondary vertex mass $m^{\text{SV}}$ became the most impactful variable in $b$ vs $c$ discrimination at high purity by a sizeable margin, as SV kinematics store essential information about the decaying hadrons.
@@ 5.5 Dependence on the Flavour Definition
Defining the flavour of a reconstructed jet is a complex task. Several definitions have been used in past and current experiments to assign the flavour of MC-generated jets, reconstructed using a detector simulation. However, flavour definitions designed for jets clustered with cone-shaped algorithms, like the anti-$k_{\text{T}}$ algorithm Cacciari_2008 , are not suitable for irregularly-shaped jets, like the ones clustered in this study with the exclusive (Durham) $e^{+}e^{-}~{}k_{\text{T}}$ algorithm Cacciari_2012 .
Various flavour definitions were considered to study their impact on the classifier performance for $Z$ boson decay events. For conciseness, a comparison of two of these definitions is reported here. The first of these is the Ghost Matching algorithm CACCIARI2008119 used at CMS, which defines the flavour of a jet by finding the hadrons or partons from the MC history of the jet, called ghosts, clustered with the jet after scaling their momentum with a minuscule factor. The other definition assigns the flavour of a jet as the flavour of the quark to which the $Z$ boson decays. A performance difference of $11.8\%$ was seen in the discrimination of $s$-jets vs $ud$-jets at a fixed background efficiency of $10\%$ between the two flavour definitions. Such a significant difference makes the considered jet flavour definition consequential while comparing different classifiers.
@@ 6 Example of Performance: The $Z$ Boson at the FCC-ee
The $Z$ boson decays relatively uniformly to the five quark flavours, and none of the decay channels to $q\bar{q}$ pairs are suppressed. Thus, tagging a particular jet flavour entails discrimination against every other flavour. Especially, isolating $Z\to s\bar{s}$ events from the exclusive decays of the $Z$ boson provides a challenging case to tag the $s$-jets by eliminating both the heavy jets and the light jets. The dominant discriminating variable against the heavy jets is the reconstructed SVs, while it is the presence of a leading strange hadron against the light jets. This makes tagging $Z\to s\bar{s}$ events an ideal metric to assess the performance of DeepJetTransformer and allows for a unique opportunity to access a hitherto scarcely studied channel.
6.1 Physics Potential at the $Z$ Resonance
Heavy-flavour tagging was performed at LEP 2004 ; Proriol:1950599 and the Tevatron Abazov_2010 ; Freeman_2013 by reconstructing SVs and exploiting these to remove the background from light-flavour jets. SLD also tagged $Z\to s\bar{s}$ events, to measure $A_{s}$, by the absence of reconstructed $B$ and $D$ hadrons and the presence of $K^{\pm}$ or $K^{0}_{S}$ PhysRevLett.85.5059 . The particle identification was performed at SLD, as at DELPHI, with a RICH detector SLD:1984aa ; delphi_loi . At most other detectors, $dE/dx$ was used for PID opal_loi ; l3_loi , with the addition of timing at ALEPH aleph_loi .
Future colliders with a dedicated $Z$ boson run, like FCC-ee, will improve the precision of all these measurements, making the lighter quarks accessible. Individual measurements of the quark vector and axial couplings should be possible via their forward-backward asymmetries, corresponding partial decay widths of the $Z$ boson, and the precise knowledge of $A_{e}$. The experimental systematic uncertainties corresponding to these measurements are also expected to drastically improve due to better detector designs with PID and vertexing Benedikt:2651299 . This section aims to evaluate the performance of DeepJetTransformer in tagging $Z\to q\bar{q}$ events. This will be demonstrated by isolating $Z\to s\bar{s}$ events from the exclusive hadronic decays of the $Z$ boson in the FCC-ee environment. Further backgrounds are not considered.
@@ 6.2 Event and Jet Selection
The simulated samples described in Sec. 2.3 are used. These samples use Pythia8.303 https://doi.org/10.48550/arxiv.2203.11601 to generate $e^{+}e^{-}\rightarrow Z\rightarrow q\bar{q}$ events, where $q\equiv b,c,(u,d,s)$, at the center-of-mass energy of $91.2$ GeV and were clustered exclusively into $2$ jets with FastJet-3.3.4 Cacciari_2012 using the $e^{+}e^{-}$ $k_{\text{T}}$ algorithm Catani:1991hj .
Events are selected if exactly two jets could be reconstructed with their final constituents. Jets with low momentum or jet axes outside the fiducial region of the detector are excluded. An event is selected if both of its jets have a momentum magnitude ($|p|$) greater than $20$ GeV and the polar angle ($\theta$) of their jet axes within $14$ and $176$ degrees. Events are required to have jets of the same MC flavour, defined as the flavour of the quarks to which the $Z$ boson decays.
@@ 6.3 Performance and Working Points
Discriminants are defined to sequentially remove the heavy flavour background ($b$ and $c$) and the light flavour background ($u$ and $d$). The $s$-jets are first tagged to be discriminated from $b$- and $c$-jets by defining the discriminant as in Eq. 4 with $s$-jets as signal and $b$- and $c$-jets as background. For the jets tagged by introducing a cut on this discriminant, another discriminant is defined to distinguish $s$-jets from $u$- and $d$-jets through the same method. The signal efficiencies after each subsequent cut, corresponding to four working points with increasing purity, are reported in Table 5.
@@
WP1 s vs bc $10.01$ $98.93\pm 0.03$ $7.35\times 10^{11}$ $1.35\times 10^{12}$
s vs ud $10.03$ $40.03\pm 0.04$ $1.45\times 10^{11}$ $3.25\times 10^{10}$
WP2 s vs bc $1.02$ $54.18\pm 0.04$ $2.38\times 10^{11}$ $2.06\times 10^{11}$
s vs ud $10.03$ $39.28\pm 0.06$ $5.10\times 10^{10}$ $5.57\times 10^{9}$
WP3 s vs bc $1.02$ $54.18\pm 0.04$ $2.38\times 10^{11}$ $2.06\times 10^{11}$
s vs ud $1.0$ $10.05\pm 0.11$ $1.12\times 10^{10}$ $4.77\times 10^{8}$
WP4 s vs bc $0.11$ $17.96\pm 0.06$ $3.23\times 10^{10}$ $6.98\times 10^{9}$
@@
Table 5: Presented are the efficiencies to select $s$ quark jets and the mistag rate for other flavours at four different working points. Also listed are the expected yields calculated for an integrated luminosity of $125~{}\text{ab}^{-1}$. Signal is defined as $Z\to s\bar{s}$ events while the background is composed of $Z\to q\bar{q}$ events (all quarks but $s$ quarks) events. The number of observed events is significantly above the canonical discovery significance of five standard deviations for all selections.
Figure 9: The reconstructed invariant mass of the dijet system before and after tagging strange jets with DeepJetTransformer, for an assumed integrated luminosity of $125~{}\mathrm{ab}^{-1}$. Shown are (a) the distribution without tagging applied, (b) after rejection of $b$- and $c$-jets vs $s$-jets at $1\%$ mistag rate, (c) the distribution after rejection of $b$- and $c$-jets at $1\%$ and $u$- and $d$-jets vs s-jets at $1\%$ mistag rate, corresponding to WP3 in Table 5.
As shown in Table 5, the working points are defined for four different sets of misidentification rates, referred to as mistag rates. Working Point 1 (WP1) corresponds to a mistag rate of $10\%$ while tagging $s$-jets versus the background of $b$- and $c$-jets and a mistag rate of $10\%$ while tagging $s$-jets versus the background of $u$- and $d$-jets. Working Point 2 (WP2) corresponds to a stricter mistag rate of $1\%$ while tagging $s$-jets versus the background of $b$- and $c$-jets, keeping the mistag rate the same as WP1 while tagging $s$-jets versus the background of $u$- and $d$-jets. Both mistag rates are $1\%$ for Working Point 3 (WP3). Working Point 4 (WP4) is the tightest scenario, with both mistag rates being $0.1\%$.
@@
These jets are first tagged to remove the background of $b$- and $c$-jets by defining the discriminant, as described above. If both jets from a $Z$ boson decay event are tagged with the same flavour, they are used to reconstruct the invariant mass. The distribution of this invariant mass after the first tag is displayed in Fig. 9(b), with the contributions of the MC flavours of the jets indicated. The events passing the anti-$b/c$ tag requirement are subsequently tagged with the $s$ vs light quark tagger to remove the background of $u$- and $d$-jets. Fig. 9(c) shows the distribution of the reconstructed invariant mass of $Z$ boson. Both jets are required to be tagged in each stage of the selection.
Figure 10: Discovery Significance vs Luminosity for all Working Points
The reconstructed tagged $Z$ resonance in Figure 9 shows that the $Z\to s\bar{s}$ sample is extremely pure after requiring two tags. Similarly, Table 5 lists events corresponding to an integrated luminosity of $125~{}\mathrm{ab}^{-1}$ that are significantly above the canonical discovery significance of $5\sigma$. It is important to realise that machine backgrounds and irreducible backgrounds from other standard model processes are not considered in this study. However, the remarkable sensitivity warrants investigation of how limited the integrated luminosity needs to be to observe $Z\to s\bar{s}$ in the considered scenario. Figure 10 shows the discovery significance of the process $Z\to s\bar{s}$, under the background-free scenario, as a function of integrated luminosity. The discovery significance, $Z$, in $\sigma$, is defined sig_calc_note as,
@@
$N_{sig}$ and $N_{bkg}$ refer to the number of signal and background events, respectively. Their corresponding values at each working point can be referred to from Table 5. It can be seen that a $5\sigma$ significance can be achieved with minuscule luminosities compared to the FCC-ee run plan, even at the tightest working point. For WP3, corresponding to Figure 9, a $5\sigma$ significance can be reached with a luminosity of $60~{}\text{nb}^{-1}$, equivalent to less than a second of the FCC-ee run at the $Z$ resonance.
These findings will open up avenues at FCC-ee for measurements that require ultra-pure $Z\to q\bar{q}$ samples, at least for the three heaviest flavours to which the $Z$ boson decays. Some examples are vector and axial couplings of the $Z$ to up- and down-type quarks and possibly even individual quark flavours and asymmetry parameters of the $Z$ boson in the hadronic decay channels. LEP and SLD performed comprehensive measurements of the forward-backwards charge asymmetry for $e^{+}e^{-}\to b\bar{b}$ Workman:2022ynf , similar precise measurements for the charm and the strange quark, and possibly the light quarks, will become feasible at the FCC-ee.

### v4

@@ 5.4 Importance of Variable Classes and Individual Variables
Aiming to estimate the relative importance of a given variable class (e.g. SV variables), the classifier performance was evaluated using the Permutation Feature Importance Breiman2001 ; fisher2019all method.
In particular, the variable class under investigation was shuffled amongst all other jets, keeping the rest of the variables unchanged. Specifically, the values for the variable class under investigation were randomly permuted across all jets in the dataset, while the remaining variables for all jets were left unchanged. This disrupts the relationship between the permuted variable and the jet classification, allowing for an estimate of how much the performance of the classifier depends on the given variable. The resulting performance change was considered for discriminating between $b$- vs $c$-, $c$- vs $s$-, and $s$- vs $ud$- jets, compared to the baseline where no variable classes were permuted. Charged jet constituent variables, listed in Table 3, were found to be the most impactful variable class for all types of discrimination at a background efficiency of $\epsilon_{bkg}=10\%$, as depicted in Table 6. This is presumably due to charged particles being the majority of the reconstructed particles in the jets. SV variables, listed in Table 4, primarily benefited $c$ vs $s$ discrimination, with $s$ vs $ud$ tagging particularly insensitive. Of the remaining three variable classes, V0 variables and neutral jet constituent variables were found to almost exclusively impact the performance of $s$ vs $ud$ discrimination, with little impact on both $b$ vs $c$ and $c$ vs $s$ discrimination, justifying the inclusion of V0s for identifying $s$-jets through conservation of strangeness. Jet-level variables were found to be the least significant, marginally impacting $s$ vs $ud$ discrimination, and having virtually no impact on heavy flavour discrimination. Moving to the high purity regime at a background efficiency of $\epsilon_{bkg}=0.1\%$, primarily the same trends were observed, with the impact of any variable type being amplified. SV variables, in particular, became hugely important to heavy flavour tagging, reaching almost equal in impact to the charged jet constituent variables, proving that the presence and properties of SVs are definitive indicators for identifying heavy flavour jets.
Variable Class Jet-level Charged Neutral SV V0
b vs c $2.4\%$ $62.4\%$ $2.2\%$ $13.9\%$ $0.1\%$
$\epsilon_{bkg}=10\%$ c vs s $1.2\%$ $65.7\%$ $2.9\%$ $29.6\%$ $0.2\%$
s vs ud $7.6\%$ $59.4\%$ $21.8\%$ $5.0\%$ $16.4\%$
b vs c $6.6\%$ $97.0\%$ $8.0\%$ $89.9\%$ $0.6\%$
$\epsilon_{bkg}=0.1\%$ c vs s $9.3\%$ $96.1\%$ $11.0\%$ $77.9\%$ $0.2\%$
s vs ud $35.9\%$ $91.0\%$ $57.3\%$ $7.4\%$ $43.8\%$
Table 6: Performance decrease in signal efficiency ($\epsilon_{sig}$) after permutation of variable classes defined in Section 4.1 for fixed background efficiencies ($\epsilon_{bkg}$) of 10% and 0.1%.
@@
  • Kinematic ($|p|$, $E$, $|p|/|p|_{\text{jet}}$, $\theta$, $\Delta\theta$, …)
@@
It was found that, at a background efficiency of $10\%$, kinematic variables of charged particle constituents, including $\dfrac{E_{\text{ch.}}}{E_{\text{jet}}}$ and $\dfrac{|p|_{\text{ch.}}}{|p|_{\text{jet}}}$, were generally impactful, particularly for $c$ vs $s$ discrimination. Track variables, such as $D_{0}/\sigma_{D_{0}}$ and $z_{0}$, were the most impactful, though less for $b$ vs $c$ than other types of discrimination, possibly due to their redundant information after the inclusion of SVs. PID variables had little impact on $b$ vs $c$ and $c$ vs $s$ discrimination, but $K^{\pm}\text{ID}$ and photon ID were the most important for $s$ vs $ud$ discrimination, as was observed earlier. The high purity regime at a background efficiency of $0.1\%$ resulted in similar trends, though with PID variables, including $K^{\pm}\text{ID}$ and photon ID, decreasing in importance and being somewhat replaced by kinematic ones. It should be stated that the baseline $K^{\pm}\text{ID}$ scenario, as mentioned in Section 5, is deliberately pessimistic, which could account for its decrease in importance. Track variables remained the most impactful. The secondary vertex mass $m^{\text{SV}}$ became the most impactful variable in $b$ vs $c$ discrimination at high purity by a sizeable margin, as SV kinematics store essential information about the decaying hadrons. The results of this study are summarised in Table 7 below.
Variable $\ln(E_{\text{ch.}})$ isPhoton $K^{\pm}\text{ID}$ $m^{\text{SV}}$ $|p|^{\text{V}^{0}}$ $z_{0}$ $D_{0}/\sigma_{D_{0}}$
b vs c $3.5\%$ $0.3\%$ $0.2\%$ $3.0\%$ $0.1\%$ $7.8\%$ $11.6\%$
$\epsilon_{bkg}=10\%$ c vs s $23.8\%$ $0.7\%$ $0.5\%$ $0.3\%$ $0.2\%$ $20.9\%$ $39.1\%$
s vs ud $12.8\%$ $16.6\%$ $38.8\%$ $0.0\%$ $9.2\%$ $23.3\%$ $26.7\%$
b vs c $13.8\%$ $1.3\%$ $0.9\%$ $67.2\%$ $0.8\%$ $34.1\%$ $45.0\%$
$\epsilon_{bkg}=0.1\%$ c vs s $57.6\%$ $0.9\%$ $4.8\%$ $7.0\%$ $0.3\%$ $56.2\%$ $79.5\%$
s vs ud $35.0\%$ $28.0\%$ $59.0\%$ $0.4\%$ $34.7\%$ $60.5\%$ $80.1\%$
Table 7: Performance decrease in signal efficiency ($\epsilon_{sig}$) after permutation of individual variables defined in Section 4.1 for fixed background efficiencies ($\epsilon_{bkg}$) of 10% and 0.1%. A set of seven variables, chosen among the most impactful, is presented here.
@@ 5.5 Dependence on the Flavour Definition
Defining the flavour of a reconstructed jet is a complex task. Several definitions have been used in past and current experiments to assign the flavour of MC-generated jets. The flavour definition can impact the classifier performance for $Z$ boson decay events because this definition can lead to jets being assigned a different flavour than the original quark.
In the $Z$ boson definition used throughout this work as introduced in Section 2.3, the flavour of a jet is defined as the flavour of the quark to which the $Z$ boson decays. The hadronisation and fragmentation of the quark are ignored in this definition. One flavour definition that accounts for fragmentation and hadronisation effects is the Ghost Matching algorithm used at CMS CACCIARI2008119 , which defines the flavour of a jet by finding the hadrons or partons from the MC history of the jet, clustered with the same jet clustering algorithm as the reconstructed jet.
The largest performance differences of the algorithm after changing flavour definitions can be observed in the discrimination of $s$-jets vs $ud$-jets, where the Ghost Matching definition leads to a 11.8% higher tagging efficiency than the $Z$ boson definition at a fixed background efficiency of 10%. Such significant changes in performance make it essential to account for the used flavour definitions while comparing different flavour tagging algorithms.
@@ 6 Example of Performance: The Z Boson at the FCC-ee
The $Z$ boson decays relatively uniformly to the five quark flavours, and none of the decay channels to $q\bar{q}$ pairs are suppressed. Thus, tagging a particular jet flavour entails discrimination against every other flavour. Especially, isolating $Z\to s\bar{s}$ events from the exclusive decays of the $Z$ boson provides a challenging case to tag the $s$-jets by eliminating both the heavy jets and $u$-, $d$-jets. The dominant discriminating variable against the heavy jets is the reconstructed SVs, while it is the presence of a high-momentum strange hadron against $u$- and $d$-jets. This makes isolating $Z\to s\bar{s}$ events from the exclusive hadronic decays of the $Z$ boson an ideal metric to assess the performance of DeepJetTransformer in the FCC-ee environment and allows for a unique opportunity to access a hitherto scarcely studied channel. Further backgrounds are not considered but are expected to be well under one per cent of the total expected yield around the $Z$ boson resonance.
@@ 6.1 Event and Jet Selection
The $e^{+}e^{-}\rightarrow Z\rightarrow q\bar{q}$ event samples with $q\equiv b,c,(u,d,s)$ described in Section 2.3 are used. These are the same samples used to evaluate the performance of DeepJetTransformer.
Events are selected if exactly two jets could be reconstructed with their final constituents. Jets with low momentum or jet axes outside the fiducial region of the detector are excluded. An event is selected if both of its jets have a 3-momentum magnitude ($|p|$) greater than $20$ GeV and the polar angle ($\theta$) of their jet axes within $14$ and $176$ degrees. Events are required to have jets of the same MC flavour, defined as the flavour of the quarks to which the $Z$ boson decays.
@@ 6.2 Performance and Working Points
All jets from $Z\to q\bar{q}$ events are independently evaluated using DeepJetTransformer. Discriminants are defined to sequentially remove the heavy flavour background ($b$- and $c$-jets) and the light flavour background ($u$- and $d$-jets). The $s$-jets are first tagged to be discriminated from $b$- and $c$-jets by defining the discriminant as in Eq. 4 with $s$-jets as signal and $b$- and $c$-jets as background. For the jets tagged by introducing a cut on this discriminant, another discriminant is defined to distinguish $s$-jets from $u$- and $d$-jets through the same method. The signal efficiencies after each subsequent cut, corresponding to four working points with increasing purity, are reported in Table 8.
@@
WP1 s vs bc $10$ $98.93\pm 0.03$ $7.35\times 10^{11}$ $1.35\times 10^{12}$
s vs ud $10$ $40.03\pm 0.04$ $1.45\times 10^{11}$ $3.25\times 10^{10}$
WP2 s vs bc $1$ $54.18\pm 0.04$ $2.38\times 10^{11}$ $2.06\times 10^{11}$
s vs ud $10$ $39.28\pm 0.06$ $5.10\times 10^{10}$ $5.57\times 10^{9}$
WP3 s vs bc $1$ $54.18\pm 0.04$ $2.38\times 10^{11}$ $2.06\times 10^{11}$
s vs ud $1$ $10.05\pm 0.11$ $1.12\times 10^{10}$ $4.77\times 10^{8}$
WP4 s vs bc $0.1$ $17.96\pm 0.06$ $3.23\times 10^{10}$ $6.98\times 10^{9}$
@@
Table 8: Presented are the efficiencies to select $s$ quark jets and the mistag rate for other flavours at four different working points. Also listed are the expected yields calculated for an integrated luminosity of $125~{}\text{ab}^{-1}$. Signal is defined as $Z\to s\bar{s}$ events while the background is composed of $Z\to q\bar{q}$ (all quarks but $s$ quarks) events. The number of observed events is significantly above the canonical discovery significance of five standard deviations for all selections.
Figure 9: The reconstructed invariant mass of the dijet system before and after tagging both jets with DeepJetTransformer, corresponding to WP2 and WP3 in Table 8, for an assumed integrated luminosity of $125~{}\mathrm{ab}^{-1}$. Both jets are required to be tagged in each case. Shown are (a) the distribution without tagging applied, (b) after the rejection of $b$- and $c$-jets vs $s$-jets at $1\%$ mistag rate, (c) the distribution after rejection of $b$- and $c$-jets at $1\%$ and $u$- and $d$-jets vs s-jets at $10\%$ mistag rate, (d) the distribution after rejection of $b$- and $c$-jets at $1\%$ and $u$- and $d$-jets vs s-jets at $1\%$ mistag rate.
@@
These jets are first tagged to remove the background of $b$- and $c$-jets by defining the discriminant, as described above. If both jets from a $Z$ boson decay event pass this tagging requirement, they are used to reconstruct the invariant mass. The distribution of this invariant mass is displayed in Figure 9(b), with the contributions of the MC flavours of the jets indicated. The jets from the events passing the anti-$b/c$ tag requirement are subsequently tagged with the $s$ vs $ud$ quark tagger to remove the background of $u$- and $d$-jets. Figures 9(c) and 9(d) show the distribution of the reconstructed invariant mass of the $Z$ boson after this additional tag. Both jets of every event are required to pass the tagging requirements in each stage of the selection.
Figure 10: Discovery Significance vs Luminosity for the four working points corresponding to Table 8. The points noted by the $x$ axis intersecting with the dashed vertical lines are the luminosities required at the four respective working points to achieve the canonical discovery significance of $5\sigma$.
The reconstructed tagged $Z$ resonance in Figure 9 shows that the $Z\to s\bar{s}$ sample is extremely pure after requiring the two consecutive tags on each jet from $Z$ boson decay events. Table 8 lists events corresponding to an integrated luminosity of $125~{}\mathrm{ab}^{-1}$. The discovery significance, $Z$, in $\sigma$, is defined sig_calc_note as,
@@
$N_{sig}$ and $N_{bkg}$ refer to the number of signal and background events, respectively. Signal is defined as $Z\to s\bar{s}$ events while the background is composed of $Z\to q\bar{q}$ (all quarks but $s$ quarks) events. It is apparent that all four working points are significantly above the canonical discovery significance of $5\sigma$. It is important to realise that machine backgrounds and irreducible backgrounds from other standard model processes are not considered in this study, and are at the per cent level. However, the remarkable sensitivity warrants investigation of how limited the integrated luminosity can be to observe $Z\to s\bar{s}$ in the considered scenario.
Figure 10 shows the discovery significance of the process $Z\to s\bar{s}$, under the background-free scenario, as a function of integrated luminosity. The corresponding values of $N_{sig}$ and $N_{bkg}$ at each working point can be referred to from Table 8. It can be seen that a $5\sigma$ significance can be achieved with minuscule luminosities compared to the FCC-ee run plan, even at the tightest working point. For WP3, corresponding to Figure 9(d), a $5\sigma$ significance can be reached with a luminosity of $60~{}\text{nb}^{-1}$, equivalent to less than a second of the FCC-ee run at the $Z$ resonance.
Data-to-simulation scale factors for $b$-jets can be measured with a precision of approximately ±$\pm 2.5\%$ for jets with $5030<p_{\text{T}}<$ 50 GeV at the LHC experiments. Tagging algorithms at the future colliders are expected to achieve smaller uncertainties.
These findings will open up avenues at FCC-ee for measurements that require ultra-pure $Z\to q\bar{q}$ samples, at least for the three heaviest flavours to which the $Z$ boson decays. Some examples are vector and axial couplings of the $Z$ to up- and down-type quarks and possibly even individual quark flavours and asymmetry parameters of the $Z$ boson in the hadronic decay channels. LEP and SLD performed comprehensive measurements of the forward-backwards charge asymmetry for $e^{+}e^{-}\to b\bar{b}$ Workman:2022ynf , similar precise measurements for the charm and the strange quark, and possibly the $u$, $d$ quarks, will become feasible at the FCC-ee.
7 Summary
The transformer-based model presented in this work can be trained considerably more quickly than the state-of-the-art graph neural network-based taggers Qu:2022mxj ; 10363595 . The discrimination power of this framework called DeepJetTransformer is presented for FCC-ee, allowing the classification of all jet flavours in $e^{+}e^{-}$ collisions at the $Z$ resonance.
A tagging efficiency for $b$-jets of about $99\%$($86\%$) can be achieved against $s$-, $u$-, and $d$-jets (c-jets) at a background efficiency of $0.1\%$, pointing to an excellent $b$-jet discrimination, dominantly owing to the secondary vertex reconstruction coming from the expected excellent detector resolution. A $c$-jet tagging efficiency of about $90\%$($70\%$) can be achieved when discriminating from $b$-jets, at a background efficiency of $10\%$($1\%$).
Excellent discrimination can be achieved for $s$-quark tagging against the $b$- and $c$-quark jet background. Against the most challenging background of $ud$-jets, a $40\%$ efficiency for $s$ quark jets can be achieved at a background efficiency of $10\%$. This performance is partially attributed to the inclusion of V0s. Further significant performance enhancement in strange tagging is seen when $K^{\pm}/\pi^{\pm}$ discrimination is included. Minor discrimination can even be achieved between $u$- and $d$-jets. The $Z\to s\bar{s}$ process can be efficiently isolated from other hadronic decays of the $Z$ boson, and an extremely pure $Z\to s\bar{s}$ sample can be obtained.